i have a test on quadratic equations, and i'm so lost~
Solve the equation algebraically. Check the
solutions graphically.

x^2-11=14

Answer by stanbon(75887)

Solve the equation algebraically. Check the
solutions graphically.
x^2-11=14
----
x^2-11-14 = 0
x^2-25 = 0
Factor:
(x-5)(x+5) = 0
x = 5 or x = -5
-------------------
Cheers,
Stan H.
========